# Accurate Identification of Communication Between Multiple Interacting Neural Populations

### v1

@@ A.2 Reduced-Rank Regression
where each $W^{\text{ridge},j}$ corresponds to the contribution of source region $A^{j}$, and $[W^{1};...;W^{N}]$ represents a vertical stack of the matrices. SVD is then applied to each individual $W^{\text{ridge},j}$ matrix:
@@ B.1 Quantifying Effectome Similarity
In a trained MR-LFADS model, we define the inferred effectome to be a matrix of pairwise message norms, $M$, with element $M_{i,j}$ as the average value of $||\mu_{m,t}^{j\rightarrow i}||_{2}$ across all trials and timesteps (see Eq. 7). In Experiments 1-2, we compared model-inferred effectomes to the corresponding ground truth connectivity matrix $M_{\text{true}}$, consisting of ones and zeros to indicate the presence or absence of a communication channel in the DGN, respectively. In contrast, Experiment 3 features networks in which not all connections are actively used; in this case, we define $M_{\text{true}}$ analogously to $M$, but computed using ground truth messages $m^{j\rightarrow i}$ instead of inferred messages $\mu_{m,t}^{j\rightarrow i}$. To assess similarity between $M$ and $M_{\text{true}}$, we flatten these matrices into vectors—$\vec{m}$ and $\vec{m}_{\text{true}}$—and compute their cosine similarity:
@@
To visualize an inferred effectome (e.g., Fig. 2c, right), we plot arrows with line thickness indicating the corresponding scalar message norm from the inferred effectome. We thresholded and clipped these scalars before plotting to improve visual clarity. The threshold is determined using $k$-means clustering on the $\||\mu_{m,t}^{j\rightarrow i}||_{2}$ values, separating them into two groups and omitting the group with smaller values. Line thickness is capped at an upper bound for readability. All reported values elsewhere are computed without applying any thresholds, and visualizations of the unthresholded effectome are also provided in Fig. S1b, Fig. S2b.
@@ C.2 Pass-Decision Network
Each region has a linear readout $W_{\text{out}}$, whose output is required to match its corresponding latent variables ($s_{t}$ for area $A^{P}$ and $d_{t}$ for $A^{D}$). The message sent from $A^{P}$ to $A^{D}$ is trained to represent $s_{t}$. Additionally, $A^{D}$ must encode whether $d_{t}$ is greater or less than 00 at all times, mimicking a binary decision-making process.
@@
Figure S1: Memory network. (a) Example traces of model reconstruction. Row: example neuron from an area, column: different trials. (b) Inferred effectomes visualized as heatmaps. Color represents message norms normalized by the largest message norm within each model, $\max_{i,j}||\mu_{m}^{j\rightarrow i}||_{2}$. These largest messages are, from left to right: $195$, $22$, $95$, $1431$, $5$. (c). $R^{2}$ of linear prediction of ground truth messages (with $2$ time step lag on the left, $4$ time step lag on the right) via inferred messages. MR-LFADS(R) results (top) are replicated from Fig. 2e.
Figure S2: Pass-decision network. (a) Example traces of model reconstruction. Row: example neuron from an area, column: different trials. (b). (b) Inferred effectomes visualized as circuit diagrams and heatmaps. Color represents message norms normalized by the largest message norm within each model, $\max_{i,j}||\mu_{m}^{j\rightarrow i}||^{2}_{2}$. These largest messages are, from left to right: $287$, $271$, $307$, $143$, $398$, $0.6$.
@@ Appendix D Randomly Generated Multi-Region Networks
We generated a distribution of networks designed to perform computational tasks inspired by Yang et al. (2019). Each network consists of either 3 or 4 regions. A connection probability $p\in\{0.5,0.6,0.7\}$ is specified, and the connectome is randomly drawn. To be considered valid, the connectome must meet two criteria: (1) each region must have at least one input connection and one output connection to ensure no region is redundant, and (2) all regions must be within a maximum distance of 2 steps from the output region. Once a valid connectome is generated, the network is trained on one of the computational tasks. During training, dynamic noise $\xi\sim\mathcal{N}(0,0.01\,\mathbb{I})$ is applied to all regions, and the only loss is based on whether the output region produces the correct response. Networks that meet the performance thresholds (train accuracy $>0.8$ and validation accuracy $>0.6$) are selected as synthetic datasets.
@@
All trials are 200 time steps in length. Each task receives a fixation input $s_{\text{fix},t}$, two stimuli $s^{\text{(1)}}_{t}=(a^{\text{(1)}},\theta^{\text{(1)}})$ and $s^{\text{(2)}}_{t}=(a^{\text{(2)}},\theta^{\text{(2)}})$, and requires a saccade response $r^{\text{sacc}}_{t}$ and an additional response $r_{\text{resp},t}=(1,\theta_{\text{resp}})$, which differ based on the specified task. Both the stimuli and responses are expressed in polar coordinates, with a resolution of 10 degrees per angle.
The tasks are described in terms of $3$ different families: the go task family, context-dependent decision-making family, and matching family. For all tasks, $t_{\text{start}}\in[30,50)$ denotes the onset of the first (and sometimes only) stimulus. The duration of all stimulus pulses in a trial is represented by $\Delta_{\text{dur}}\in[30,50)$, while $\Delta_{\text{offset}}$ specifies the offset between the two stimuli, if applicable. The time between the last stimulus offset and the fixation cue offset is given by $\Delta_{\text{delay}}$, where $\Delta_{\text{delay}}=\infty$ indicates that the fixation cue never disappears. The onset of the saccade is denoted as $t_{\text{sacc}}$.
@@ D.3 Matching Family
Figure S3: Randomly generated multi-region network: example traces of model reconstruction for networks with median MR-LFADS(R) performance for 3 regions (a) and 4 regions (b).
@@ E.1 Input and Message Inference with KL Penalties
Figure S4: The effect of KL penalties on restricting information across inferred inputs and communication channels in Experiment 1. (a) Inferred input norm over time, $||\mu_{u,t}^{[k]}||_{2}$, for the most (red) and least (pink) active channels, where $k$ denotes channel number. (b) Inferred message norm over time, $||\mu_{m,t}^{j\rightarrow i,[k]}||_{2}$, for the most (blue) and least (cyan) active channels. (c) Inferred input / message norm for area $A^{1}$.
@@ E.2 Message Inference via Rates versus Factors
Since MR-LFADS factors and reconstructed rates share the same dimensionality and are related by a linear transformation, it may not be immediately clear how message inference from these variables leads to substantial differences. To investigate this, we examine MR-LFADS(F) trained on the pass-decision synthetic data. We perform SVD on projection matrices from factors to rates, $W_{r}$, and from factors to communication, $W_{m}$ (Fig. S5a). Our analysis shows that $W_{r}$ has small singular values, indicating that certain factor dimensions contribute minimally to the reconstructed rates (Fig. S5b). In contrast, $W_{m}$ exhibits relatively uniform singular values (within the same order of magnitude), suggesting that all factor dimensions are utilized in communication (Fig. S5c). This implies that some factor dimensions play a role in message inference while being largely detached from rate reconstruction (Fig. S5a).
To further investigate, we re-express the factor space using the left singular vectors of $W_{r}$, denoted $U$. In Experiment 2, MR-LFADS(F) is shown to encode both the stimulus $s$ and decision variable $d$ in its $m^{P\rightarrow D}$ messages (Fig. 3e). To examine how these variables are distributed across the subspaces of $U$ and whether this aligns with the under-constrained dimensions identified earlier, we project the factors $f$ onto the subspace spanned by the top $k$ singular vectors of $U$, denoted $U^{[1:k]}$, varying $k$ from 1 to $N_{\text{msg}}^{D}$. We then compute the $R^{2}$ values for predicting $s$ and $d$ from the projected values $f^{\prime}$ (Fig. S5d, e, cyan lines). We repeat this process in the reverse direction, projecting onto the bottom $k$ singular vectors instead (Fig. S5d, e, orange lines). The results reveal a clear separation: decoding accuracy for $s$ improves more when projecting onto the top singular vectors, whereas decoding accuracy for $d$ increases more rapidly when projecting onto the bottom singular vectors. This suggests that information not used for rate reconstruction—such as $d$—is preferentially encoded in the under-constrained dimensions of the latent space, reinforcing the idea that message inference utilizes latent dimensions beyond those needed for rate prediction.
Figure S5: Potential pitfall of overexpressive models. (a) Illustration of factors containing unconstrained information. $k$ is the effective rank of $W_{r}$. (b) Ranked singular values for $W_{r}$. (c) Ranked singular values for $W_{m}$. (d) $R^{2}$ for decoding $s$ from $f^{P}$ projected onto subspaces spanned by top / bottom left singular vectors of $W_{r}$ of an example seed. (e) Same as (d), but for decoding $d$.
Appendix F Application to Large Scale Electrophysiological Data
@@ F.1 Comparison of Photoinhibition Effects
We selected one session of the data involving five brain regions previously implicated in a decision-making task, as identified in Chen et al. (2024). For both control and photoinhibition trials, we only used trials from the same condition per comparison—either left-hit or right-hit—where “hit” indicates a correct choice, and “left” or “right” refers to the correct decision side. For photoinhibition trials, we focused on those with perturbations within the delay period ($\sim 2$ s), aligning all such trials to the photoinhibition onset. Firing rate $\bar{r}_{t}$ is smoothed from raw spike counts using a causal exponential filter with rate parameter of $7.1$. For each region, we computed the absolute difference in trial-averaged activity between photoinhibited and control trials, averaged over neurons, to estimate the influence of photoinhibition.

### v2

@@ A.2 Reduced-Rank Regression
where each $W^{\text{ridge},j}$ corresponds to the contribution of source region $A^{j}$, and $W^{1};...;W^{N}]$ represents a vertical stack of the matrices. SVD is then applied to each individual $W^{\text{ridge},j}$ matrix:
@@ B.1 Quantifying Effectome Similarity
[…]
@@
To visualize an inferred effectome (e.g., Fig. 2c, right), we plot arrows whose color intensity reflects the relative message norm, computed by concatenating all multi-dimensional messages across trial and time, taking the 2-norm, and then normalizing across all channels, with inferred communication and inputs normalized separately. The arrows in Fig. 4e, bottom are further scaled by a sigmoid function for visual contrast. All reported values elsewhere are computed without applying any thresholds or scaling, and heatmap visualizations of the effectome are also provided in Fig. S1b, Fig. S2b.
@@ C.2 Pass-Decision Network
Each region has a linear readout $W_{\text{out}}$, whose output is required to match its corresponding latent variables ($s_{t}$ for area $A^{P}$ and $d_{t}$ for $A^{D}$). The message sent from $A^{P}$ to $A^{D}$ is trained to represent $s_{t}$. Additionally, $A^{D}$ must encode whether $d_{t}$ is greater or less than 0 at all times, mimicking a binary decision-making process.
@@
[…]
Figure S2: Pass-decision network. (a) Example traces of data reconstruction. Row: example neuron from an area, column: different trials. (b) Inferred effectomes visualized as circuit diagrams and heatmaps. Top: circuit diagram, middle: inferred inputs, bottom: communication. Color intensity represents message norms normalized by the largest input or message norm within each model.
@@ Appendix D Randomly Generated Multi-Region Networks
We generated a distribution of networks designed to perform computational tasks inspired by Yang et al. (2019). Each network consists of either 3 or 4 regions. A connection probability $p\in\{0.5,0.6,0.7\}$ is specified, and the connectome is randomly drawn. To be considered valid, the connectome must meet two criteria: (1) each region must have at least one input connection and one output connection to ensure no region is redundant, and (2) all regions must be within a maximum distance of 2 steps from the output region. Once a valid connectome is generated, the network is trained on one of the computational tasks. During training, dynamic noise $\xi\sim\mathcal{N}(0,0.01\,\mathbb{I})$ is applied to all regions, and the only loss is based on whether the output region produces the correct response. Networks that meet the performance thresholds (train accuracy >0.8>0.8 and validation accuracy >0.6>0.6) are selected as synthetic datasets.
@@
All trials are 200 time steps in length. Each task receives a fixation input $s_{\text{fix},t}$, stimuli from two channels $s^{\text{(1)}}_{t}=(a^{\text{(1)}},\theta^{\text{(1)}})$ and $s^{\text{(2)}}_{t}=(a^{\text{(2)}},\theta^{\text{(2)}})$, and requires a saccade response $r^{\text{sacc}}_{t}$ and an additional response $r_{\text{resp},t}=(1,\theta_{\text{resp}})$, which differ based on the specified task. For different tasks, stimuli may come from one or both of the channels. Both the stimuli and responses are expressed in polar coordinates, with a resolution of 10 degrees per angle.
[…]
@@ D.3 Matching Family
Figure S3: Randomly generated multi-region network: example traces of data reconstruction for networks with median MR-LFADS(R) performance for 3 regions (a) and 4 regions (b).
@@ E.1 Input and Message Inference with KL Penalties
Figure S4: The effect of KL penalties on restricting information across inferred inputs and communication channels in Experiment 1. (a) Inferred input norm over time, $\mu_{u,t}^{[k]}||_{2}$, for the most (dark red) and least (light red) active channels, where $k$ denotes channel number. (b) Inferred message norm over time, $\mu_{m,t}^{j\rightarrow i,[k]}||_{2}$, for the most (dark blue) and least (light blue) active channels. (c) Inferred input (top) and message (bottom) norm across channels for area $A^{1}$ in descending order of its scalar norm across trials and time.
@@ E.2 Message Inference via Rates versus Factors
Since MR-LFADS factors and reconstructed rates share the same dimensionality and are related by a linear transformation, it may not be immediately clear how message inference from these variables leads to substantial differences. To investigate this, we examine MR-LFADS(F) trained on the pass-decision synthetic data. We perform SVD on projection matrices (Fig. S5a) from factors to rates, $W_{r}$ (Eq. 5), and from factors to communication, $W_{\mu_{m}}$ (Eq. 8 with $r^{j}_{t}$ replaced by $f^{j}_{t}$). Our analysis shows that $W_{r}$ has small singular values, indicating that certain factor dimensions contribute minimally to the reconstructed rates (Fig. S5b). In contrast, $W_{\mu_{m}}$ exhibits relatively uniform singular values (within the same order of magnitude), suggesting that all factor dimensions are utilized in communication (Fig. S5c). This implies that some factor dimensions play a role in message inference while being largely detached from rate reconstruction (Fig. S5a).
To further investigate, we re-express the factor space of area P using the left singular vectors of $W_{r}$, denoted $U$. In Experiment 2, MR-LFADS(F) is shown to encode both the stimulus $s$ and decision variable $d$ in its $m^{P\rightarrow D}$ messages (Fig. 3e). To examine how these variables are distributed across the subspaces of $U$ and whether this aligns with the under-constrained dimensions identified earlier, we project the factors $f^{P}$ onto the subspace spanned by the top $k$ singular vectors of $U$, denoted $U^{[1:k]}$, varying $k$ from 1 to $N_{\text{msg}}^{D}$. We then compute the $R^{2}$ values for predicting $s$ and $d$ from the projected values $f^{\prime}$ (Fig. S5d, e, cyan lines). We repeat this process in the reverse direction, projecting onto the bottom $k$ singular vectors instead (Fig. S5d, e, orange lines). The results reveal a clear separation: decoding accuracy for $s$ improves more when projecting onto the top singular vectors, whereas decoding accuracy for $d$ increases more rapidly when projecting onto the bottom singular vectors. This suggests that information not used for rate reconstruction—such as dd—is preferentially encoded in the under-constrained dimensions of the latent space, reinforcing the idea that message inference utilizes latent dimensions beyond those needed for rate prediction.
Figure S5: SVD Analysis on MR-LFADS(F). (a) Illustration of factors containing unconstrained information. Rates, factors, and communication are all from area P, i.e. $r\equiv r^{p}$, $f\equiv f^{p}$, $\mu_{m}\equiv\mu_{m}^{P\rightarrow D}$. $k$ is the effective rank of $W_{r}$. (b) Ranked singular values for $W_{r}$. (c) Ranked singular values for $W_{\mu_{m}}$. Shaded regions indicate standard deviation across different random initializations (seeds). (d) $R^{2}$ value for decoding $s$ from $f^{P}$ projected onto subspaces spanned by top / bottom left singular vectors of $W_{r}$ of an example seed. (e) Same as (d), but for decoding $d$.
Appendix F Application to Multi-Region Electrophysiological Data
@@ F.1 Comparison of Photoinhibition Effects
We selected one session of the data involving five brain regions previously implicated in a decision-making task, as identified in Chen et al. (2024). For both control and photoinhibition trials, we only used trials from the same condition per comparison—either left-hit or right-hit—where “hit” indicates making a correct choice, and “left” or “right” refers to the correct choice. For photoinhibition trials, we focused on those with perturbations within the delay period, aligning all such trials to the photoinhibition onset. Firing rate $\bar{r}_{t}$ is smoothed from raw spike counts using a causal exponential filter with rate parameter of $7.1$. For each region, we computed the absolute difference in trial-averaged activity between photoinhibited and control trials, averaged over neurons, to estimate the influence of photoinhibition.